# Fractal ground state of ion chains in periodic potentials

### v1

@@ Acknowledgments
The authors acknowledges helpful discussions with and comments of M.-C. Banuls, C. Cormick, E. Demler, S. B. Jäger, H. Landa, and V. Stojanovic, as well as the contribution of A. A. Buchheit in the preliminary phase of this project. G.M. is deeply indebted to C. Bassi-Angelini and E. Auerbach for inspiring comments. R.M. thanks B. Pascal for her precious insight. This work has been supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), with the CRC-TRR 306 “QuCoLiMa”, Project-ID No. 429529648, and by the German Ministry of Education and Research (BMBF) via the project NiQ (”Noise in Quantum Algorithm”) and via the QuantERA project NAQUAS. Project NAQUAS has received funding from the QuantERA ERA-NET Cofund in Quantum Technologies implemented within the European Union’s Horizon 2020 program. J.Y.M. was supported by the European Social Fund REACT EU through the Italian national program PON 2014-2020, DM MUR 1062/2021. M.L.C. acknowledges support from the MIT-UNIPI program, by the National Quantum Science and Technology Institute (NQSTI), spoke 10, funded under the National Recovery and Resilience Plan (NRRP), Mission 4 Component 2 Investment 1.3 - Call for tender No. 341 of 15/03/2022 of Italian Ministry of University and Research, funded by the European Union NextGenerationEU, award number PE0000023, Concession Decree No. 1564 of 11/10/2022 adopted by the Italian Ministry of University and Research, CUP D93C22000940001, and by the project PRA_2022_2023_98 ”IMAGINATION” from the University of Pisa. VV acknowledges support from the NSF Physics Frontiers Center (PHY-2317134 ) and NSF grant PHY-2207996. This research was supported in part by grants NSF PHY-1748958 and PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).

### v2

@@ Acknowledgments
The authors acknowledges helpful discussions with and comments of M.-C. Banuls, C. Cormick, E. Demler, S. B. Jäger, H. Landa, and V. Stojanovic, as well as the contribution of A. A. Buchheit in the preliminary phase of this project. G.M. is deeply indebted to C. Bassi-Angelini and E. Auerbach for inspiring comments. R.M. thanks B. Pascal for her precious insight. This work has been supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), with the CRC-TRR 306 ”QuCoLiMa”, Project-ID No. 429529648, and by the German Ministry of Education and Research (BMBF) via the project NiQ (”Noise in Quantum Algorithm”) and via the QuantERA project NAQUAS. Project NAQUAS has received funding from the QuantERA ERA-NET Cofund in Quantum Technologies implemented within the European Union’s Horizon 2020 program. J.Y.M. was supported by the European Social Fund REACT EU through the Italian national program PON 2014-2020, DM MUR 1062/2021. M.L.C. acknowledges support from the National Centre on HPC, Big Data and Quantum Computing - SPOKE 10 (Quantum Computing) and received funding from the European Union Next-GenerationEU - National Recovery and Resilience Plan (NRRP) – MISSION 4 COMPONENT 2, INVESTMENT N. 1.4 – CUP N. I53C22000690001. M.L.C. acknowledge support from the project PRA_2022_2023_98 ”IMAGINATION” and acknowledges support from the MIT-UNIPI program. V.V. acknowledges support from the NSF Physics Frontiers Center (PHY-2317134 ) and NSF grant PHY-2207996. This research was supported in part by grants NSF PHY-1748958 and PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).